# Goal-Oriented Status Updating for Real-time Remote Inference over Networks with Two-Way Delay

### v3

@@ Abstract
We study a setting where an intelligent model (e.g., a pre-trained neural network) predicts the real-time value of a target signal using data samples transmitted from a remote source according to a scheduling policy. The scheduler decides on i) the age of the samples to be sent, ii) when to send them, and iii) the length of each packet (i.e., the number of samples contained in each packet). The dependence of inference quality on the Age of Information (AoI) for a given packet length is modeled by a general relationship. Previous work assumed i.i.d. transmission delays with immediate feedback or were restricted to the case where inference performance degrades as the input data ages. Our formulation, in addition to capturing non-monotone age dependence, also covers Markovian delay on both forward and feedback links. We model this as an infinite-horizon average-cost Semi-Markov Decision Process. We obtain a closed-form solution that decides on (i) and (ii) for any constant packet length. The solution for when to send is an index-based threshold policy, where the index function is expressed in terms of the delay state and AoI at the receiver. The age of the packet selected is a function of the delay state. We separately optimize the value of the constant length. We also develop an index-based threshold policy for the variable length case, which allows a complexity reduction. In simulation results, we observe that our goal-oriented scheduler drops inference error down to one sixth with respect to age-based scheduling of unit-length packets.
Index Terms:
Remote Inference, Goal-oriented Communication, Delay with Memory, Markovian Delay, Two-way Delay, Age of Information, and Dynamic Programming.
@@ I Introduction
Today, many previously complex computational problems enjoy the aid of AI. Among these AI-aided approaches are intelligent models, e.g., neural networks. These models have played a key role in enabling certain technologies as digital twins, industrial robotics, and self-driving cars. However, these models rely on data to be delivered to them, which in the remote computation context, depends on the network delivering the relevant data in a timely fashion. This is especially interesting in the case of unavoidable and time-varying transmission delays and is within the realm of Goal-oriented Networking for AI [2].
@@
An extreme example of monitoring over long-distance network connections is a spacecraft digital twin. Consider training a digital twin on Earth based on data coming from rovers on Mars or calibrating the virtual replica of the spacecraft/satellite using live telemetry data [3, 4]. Furthermore, in recent years, digital twins have also gained importance in the monitoring of processes in industrial environments, especially in hazardous or hard-to-reach environments or when it is impractical to be on site for monitoring. [^1] A screenshot of the digital twin of the METU high voltage testing laboratory is available at https://cng-eee.metu.edu.tr/wp-content/Digital_twin.pdf?_t=1726991660. The digital twin can be monitored over the Internet, and device settings in the physical experiment configuration on site can be remotely changed on demand directly on the twin.
@@
We focus on goal-oriented communication design for real-time estimation and inference performed by intelligent models [5, 6], which are often essential for such applications. Goal-oriented communication has been recently introduced to address the scalability problem in next-generation communication networks. This concept emphasizes that networks can facilitate the effective accomplishment of the task at the destination rather than merely addressing the transmission problem [7], which aims to reliably transmit data produced by a source. Solving the effective communication problem with efficient use of network resources requires combining the data generation and transmission processes, ensuring that data samples most significant to the computation at the destination are delivered in a timely manner [8]. This implies that the communication link should select which samples to transmit based on the state of the network (e.g., the delay state). Recent efforts in the communication and control communities highlight the value of this approach in reducing communication requirements while maintaining application performance [9, 10, 11, 12, 13, 14].
@@
Age of Information (AoI) at time $t$, denoted by $\Delta(t)$, is defined in [15] as $\Delta(t)=t-U_{t}$, where $U_{t}$ is the generation time of the most recently delivered data packet. The application-layer real-time estimation and inference performance is mapped as a function of the AoI. This function is all that is required for the link layer to operate in a goal-oriented manner. This work builds upon the findings presented in a series of papers [5, 16, 6, 17, 1, 18, 19, 20], which first demonstrated the effectiveness of AoI function as a surrogate metric for enhancing timely communication for remote estimation and inference.
@@
In this paper, we study a remote inference problem, where an intelligent, pre-trained model at the receiver is utilized to infer the real-time value of a target signal, $Y_{t}$, based on a data packet, $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$, transmitted from a remote location. The data packet $X_{t-\delta}^{l}$ comprises samples of the source signal $V_{t}$ at the transmitter. Here, $\delta$ represents the AoI of the freshest sample in the packet, and $l$ denotes the packet length, corresponding to the number of samples included in the packet.
@@
Such non-monotonicity between the inference error and AoI arises when the source signal, $V_{t}$, and the target signal, $Y_{t}$, exhibit a delayed relationship, such as $Y_{t}=f(V_{t-\Upsilon})$ [6, 22, 20]. For instance, $\Upsilon$ may represent a combination of the communication delay between the controller and the actuator and the actuation delay in a physical networked control system. An example of this scenario is illustrated in the leader-follower robot setup in [6, Fig. 2]. Additionally, non-monotonicity also occurs when the relationship between $V_{t}$ and $Y_{t}$ exhibits periodicity, as shown in the temperature prediction example in [6, Fig. 4]. In such cases, when communication delays prevent the transmission of sufficiently fresh samples to the receiver, transmitting aged samples can exploit the periodicity to enhance performance.
Moreover, the inference error is non-increasing with packet length for a given AoI [17], as additional information can only improve the estimation accuracy. However, this improvement comes at the cost of longer transmission delays. In our communication design, we investigate the interplay between packet length and transmission delay, as well as the possibly non-monotonic relationship between inference error and AoI.
@@
  • We extend the system models in [6, 17] to more practical scenarios by considering Markovian delay on both forward and feedback links. Based on this extension, we formulate and solve a learning and communication co-design problem for real-time remote inference. We consider both time-invariant and time-variable packet length selection to address the varying computational capabilities of practical systems. The derived optimal scheduling policies minimize the time-average inference error for a given, possibly non-monotonic, inference error function corresponding to a particular remote inference application, thereby making the communication goal-oriented.
  • We model the learning and communication co-design problem under time-invariant packet length selection as an infinite-horizon average-cost Semi-Markov Decision Process (SMDP). Such problems are often solved using dynamic programming [23, 24] and typically do not have closed-form solutions. However, we derive a closed-form solution to this problem for any constant packet length, presented in Theorem 1. The solution consists of two main parts: (i) the AoI of data packets to be sent are determined as a stationary function of the network state, i.e., the delay state, and (ii) the transmission instants are governed by an index-based threshold policy. The index function is expressed in terms of the delay state of the network and the AoI at the receiver. We then separately optimize the value of constant packet length.
  • We also model the learning and communication co-design problem under time-variable packet length selection as an infinite-horizon average-cost SMDP. We then formulate a Bellman optimality equation and derive a structural result regarding the optimal solution. Motivated by this result, we present a simplified version of the Bellman optimality equation in Theorem 2. Solving the simplified Bellman equation using dynamic programming has significantly lower time complexity compared to solving the original Bellman optimality equation.
  • We conduct two experiments to evaluate the performance of our optimal scheduling policies: (i) remote inference of auto-regressive (AR) processes, offering a model-based evaluation; and (ii) cart-pole state prediction, providing a trace-driven evaluation. In simulation results, we observe that our goal-oriented scheduler drops inference error down to one sixth with respect to age-based scheduling of unit-length packets.
@@ I-A Related Work
The concept of Age of Information (AoI) has attracted significant research interest; see, e.g., [22, 15, 18, 21, 25, 19, 26, 27, 28, 29, 30, 31, 32, 17, 20, 6, 5, 33, 34] and a recent survey [35]. Initially, research efforts were centered on analyzing and optimizing the average AoI and peak AoI in communication networks [15, 18, 21, 25]. Recent research endeavors have revealed that the performance of real-time applications can be modeled as non-linear functions of AoI, leading to the study of optimizing these non-linear functions in control system scenarios [28, 36], remote estimation [16, 26, 31], and remote inference [22, 5, 17, 6]. While a number of studies have analyzed AoI in queuing models, closest to the spirit of this paper is the control of AoI via replacement of exogenous data arrivals with the generation of data ”at will” [37, 18, 38, 39, 19, 27, 28, 29]. A generalization of this approach is to incorporate jointly optimal sampling and scheduling policies to control not only AoI but a more sophisticated end-to-end distortion criterion by using AoI as an auxiliary parameter [16, 26, 33, 17, 5, 6]. While strikingly more demanding of analysis, these formulations take us closer to goal-oriented communication design.
@@ II System Model
We consider a time-slotted remote inference system, which comprises a data source, a transmitter, and a receiver, as illustrated in Fig. 1. On the transmitter side, the source signal $V_{t}$ is regularly sampled in each time slot $t$. The samples are stored in a buffer containing the most recent $B$ samples $\big{(}V_{t},V_{t-1},\ldots,V_{t-B+1}\big{)}$, waiting to be sent to the receiver. On the receiver side, a predictor, e.g., a pre-trained neural network, is used to infer the real-time value of a target signal $Y_{t}$ of interest. In practical time-series forecasting algorithms, the predictor needs to use multiple consecutive samples, e.g., $X_{t-b}^{l}=(V_{t-b},V_{t-b-1},\ldots,V_{t-b-l+1})$, to infer $Y_{t}$. We will design a transmission scheduler that determines (i) when to send a group of consecutive samples over the communication network to the receiver and (ii) which group of consecutive samples in the buffer to send. This medium access model is called “selection-from-buffer,” which was proposed recently in [5, 6]. Such remote inference systems are crucial for numerous real-time applications, including sensor networks, airplane/vehicular control, robotics networks, and digital twins where timely and accurate remote monitoring and control are essential for maintaining system stability and performance.
@@
We assume that the system starts operation at time slot $t=0$. In the selection-from-buffer model, a new sample of the source signal $V_{t}\in\mathcal{V}$ is added to the buffer at the beginning of each time slot $t$, and the oldest sample is discarded. The buffer, with size $B$, is assumed to be initially full, storing the samples $\big{(}V_{0},V_{-1},\ldots,V_{-B+1}\big{)}$. Thus, at any time slot $t$, the buffer contains the most recent $B$ samples $\big{(}V_{t},V_{t-1},\ldots,V_{t-B+1}\big{)}$. The transmitter sends status update packets to the receiver one by one. Let $S_{i}$ and $D_{i}$ represent the transmission start time and delivery time of the $i$-th transmitted packet, respectively, such that $S_{i}<D_{i}$. At time slot $t=S_{i}$, the scheduler selects a group of consecutive samples from the buffer and forms the packet $X_{S_{i}-b_{i}}^{l_{i}}=\big{(}V_{S_{i}-b_{i}},V_{S_{i}-b_{i}-1},\ldots,V_{S_{i}-b_{i}-l_{i}+1}\big{)}\in\mathcal{V}^{l_{i}}$. Here, $l_{i}\in\{1,\ldots,B\}$ represents the sequence length of the $i$-th packet $X_{S_{i}-b_{i}}^{l_{i}}$ while $b_{i}$ represents the relative position of the sample $V_{S_{i}-b_{i}}$ in the buffer. When $B=1$, the selection-from-buffer model reduces to the generative-at-will model which has been studied extensively in previous works [37, 18, 38, 39, 19, 27, 28, 29]. Hence, the packet $X_{S_{i}-b_{i}}^{l_{i}}$ is submitted to the communication network at time slot $S_{i}$ and delivered to the receiver at time slot $D_{i}$. Upon delivery, the receiver sends an acknowledgment (ACK) back to the transmitter, which is received at time slot $A_{i}$. We assume that the transmitter always waits for feedback before submitting a new packet. In other words, the transmitter remains silent during the time slots between $D_{i}$ and $A_{i}$ such that $S_{i}<D_{i}<A_{i}\leq S_{i+1}$ for all $i$. Fig. 2 provides an illustration of the $i$-th and the $(i+1)$-th packet transmission epochs.

### v4

@@ Abstract
We study a setting where an intelligent model (e.g., a pre-trained neural network) infers the real-time value of a target signal using data samples transmitted from a remote source. The transmission scheduler decides (i) the freshness of packets, (ii) their length (i.e., the number of samples they contain), and (iii) when they should be transmitted. The freshness is quantified using the Age of Information (AoI), and the inference quality for a given packet length is a general function of AoI. Previous works assumed i.i.d. transmission delays with immediate feedback or were restricted to the case where inference performance degrades as the input data ages. Our formulation, in addition to capturing non-monotone age dependence, also covers Markovian delay on both forward and feedback links. We model this as an infinite-horizon average-cost Semi-Markov Decision Process. We obtain a closed-form solution that decides on (i) and (iii) for any constant packet length. The solution for when to transmit is an index-based threshold policy, where the index function is expressed in terms of the delay state and AoI at the receiver. In contrast, the freshness of the selected packet is a function of only the delay state. We then separately optimize the value of the constant packet length. Moreover, we also develop an index-based threshold policy for the time-variable packet length case, which allows a complexity reduction. In simulation results, we observe that our goal-oriented scheduler drops inference error down to one-sixth with respect to the age-based scheduling of unit-length packets.
@@ I Introduction
Today, many previously complex computational problems enjoy the aid of AI. Among these AI-aided approaches are intelligent models, e.g., neural networks. These models have played a key role in enabling certain technologies as digital twins, industrial robotics, and self-driving cars. However, these models rely on data to be delivered to them, which, in the remote computation context, depends on the network delivering the relevant data in a timely fashion. This is especially interesting in the case of unavoidable and time-varying transmission delays and is within the realm of Goal-oriented Networking for AI [2, 3].
@@
An extreme example of monitoring over long-distance network connections is a spacecraft digital twin. Consider training a digital twin on Earth based on data coming from rovers on Mars or calibrating the virtual replica of the spacecraft/satellite using live telemetry data [4, 5]. Furthermore, in recent years, digital twins have also gained importance in the monitoring of processes in industrial environments, especially in hazardous or hard-to-reach environments or when it is impractical to be on site for monitoring. [^1] A screenshot of the digital twin of the METU high voltage testing laboratory is available at https://cng-eee.metu.edu.tr/wp-content/Digital_twin.pdf?_t=1726991660. The digital twin can be monitored over the Internet, and device settings in the physical experiment configuration on site can be remotely changed on demand directly on the twin. In this paper, we focus on goal-oriented communication design for real-time estimation and inference performed by intelligent models [6, 7, 8], which are often essential for such applications.
@@
Goal-oriented communication has been recently introduced to address the scalability problem in next-generation massive machine-type communication networks. This concept emphasizes that networks can facilitate the effective accomplishment of tasks at the destination rather than merely addressing the transmission problem [9], which aims to reliably transmit data produced by a source. Solving the effective communication problem with efficient use of network resources requires combining the data generation and transmission processes, ensuring that the samples most relevant to the computation at the destination are delivered in a timely manner [2]. This implies that the communication link should select which samples to transmit based on the state of the network (e.g., the delay state). Recent efforts in the communication and control communities highlight the value of this approach in reducing communication requirements while maintaining application performance [10, 11, 12, 13, 14, 15].
@@
Age of Information (AoI) at time $t$, denoted by $\Delta(t)$, is defined in [16] as $\Delta(t)=t-U_{t}$, where $U_{t}$ is the generation time of the most recently delivered data packet. The application-layer real-time estimation and inference performance is mapped as a function of the AoI. This function is all that is required for the link layer to operate in a goal-oriented manner. This work builds upon the findings presented in a series of papers [6, 17, 7, 8, 18, 19, 20], which first demonstrated the effectiveness of AoI function as a surrogate metric for enhancing timely communication for remote estimation and inference.
@@
In this paper, we study a remote inference problem, where an intelligent, pre-trained model at the receiver is utilized to infer the real-time value of a target signal, $Y_{t}$. This inference is based on the most recently received data packet, denoted as $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$}, which has been transmitted from a remote location. The data packet $X_{t-\delta}^{l}$ comprises samples of the source signal $V_{t}$ at the transmitter. Here, $\delta$ represents the AoI of the freshest sample in the packet, and $l$ denotes the packet length, corresponding to the number of samples included in the packet.
@@
Such non-monotonicity between the inference error and AoI arises when the source signal, $V_{t}$, and the target signal, $Y_{t}$, exhibit a delayed relationship, such as $Y_{t}=f(V_{t-\Upsilon})$ [7, 22, 20]. For instance, $\Upsilon$ may represent a combination of the communication delay between the controller and the actuator and the actuation delay in a physical networked control system. An example of this scenario is illustrated in the leader-follower robot setup in [7, Fig. 2]. Additionally, non-monotonicity also occurs when the relationship between $V_{t}$ and $Y_{t}$ exhibits periodicity, as shown in the temperature prediction example in [7, Fig. 4]. In such cases, when communication delays prevent the transmission of sufficiently fresh samples to the receiver, transmitting aged samples can exploit the periodicity to enhance performance.
The inference error is non-increasing with packet length for a given AoI [8], as additional information can only improve the estimation accuracy. However, this improvement comes at the cost of longer transmission delays. In our communication design, we investigate the interplay between packet length and transmission delay, along with the possibly non-monotonic relationship between inference error and AoI.
@@
  • We extend the system models in [7, 8] to more practical scenarios by considering Markovian delay on both forward and feedback links. Based on this extension, we formulate and solve a learning and communication co-design problem for real-time remote inference. We consider both time-invariant and time-variable packet length selection to address the varying computational capabilities of practical systems. The derived optimal packet selection and scheduling policies minimize the time-average inference error for a given, possibly non-monotonic, inference error function corresponding to a particular remote inference application, thereby making the communication goal-oriented.
  • We formulate the learning and communication co-design problem under time-invariant packet length selection as a two-layer nested optimization problem. The inner layer determines the freshness and transmission times of data packets for a given constant packet length $l$ and is modeled as an infinite-horizon average-cost Semi-Markov Decision Process (SMDP). Such problems are often solved using dynamic programming [23, 24] and typically do not have closed-form solutions. However, we derive a closed-form solution to this inner layer problem, presented in Theorem 1. The solution consists of two main parts: (i) the freshness of the data packets to be transmitted is determined by a stationary function of the network state, i.e., the delay state, in the previous epoch, and (ii) the transmission instants are controlled by an index-based threshold policy. The index function is expressed in terms of the delay state of the network in the previous epoch and the AoI at the receiver. We then separately optimize the value of the constant packet length $l$ in the outer layer.
  • We model the learning and communication co-design problem under time-variable packet length selection directly as a single infinite-horizon average-cost SMDP. We then formulate a Bellman optimality equation and derive a structural result regarding the optimal solution. Motivated by this result, we present a simplified version of the Bellman optimality equation in Theorem 2. Solving the simplified Bellman equation using dynamic programming has significantly lower time complexity compared to solving the original Bellman optimality equation.
  • We conduct two experiments to evaluate the performance of our optimal scheduling policies: (i) remote inference of auto-regressive (AR) processes, offering a model-based evaluation; and (ii) cart-pole state prediction, providing a trace-driven evaluation. In simulation results, we observe that our goal-oriented scheduler drops inference error down to one-sixth with respect to the age-based scheduling of unit-length packets.
@@ I-A Related Work
The concept of Age of Information (AoI) has attracted significant research interest; see, e.g., [22, 16, 18, 21, 25, 19, 26, 27, 28, 29, 30, 31, 32, 8, 20, 7, 6, 33, 34] and a recent survey [35]. Initially, research efforts were centered on analyzing and optimizing the average AoI and peak AoI in communication networks [16, 18, 21, 25]. Recent research endeavors have revealed that the performance of real-time applications can be modeled as non-linear functions of AoI, leading to the study of optimizing these non-linear functions in control system scenarios [28, 36], remote estimation [17, 26, 31], and remote inference [22, 6, 8, 7].
While many studies have analyzed AoI in queuing models, closest to the spirit of this paper is the control of AoI via replacement of exogenous data arrivals with the generation of data ”at will” [37, 18, 38, 39, 19, 27, 28, 29]. A generalization of this approach is to incorporate jointly optimal sampling and scheduling policies to control not only AoI but a more sophisticated end-to-end distortion criterion by using AoI as an auxiliary parameter [17, 26, 33, 8, 6, 7]. While strikingly more demanding of analysis, these formulations take us closer to goal-oriented communication design.
@@
The earlier version of this paper [1] is a special case of the learning and communication co-design problem under time-invariant packet length selection discussed in Section III. In [1], the scheduler is restricted to transmitting a single sample at each transmission, i.e., the packet length is fixed to $l=1$. In contrast, the current paper allows the scheduler to form packets comprising multiple consecutive samples for transmission. Packets containing a larger number of samples can yield better inference performance. However, the associated increase in transmission delay makes such packets staler upon arrival, which may eventually degrade inference performance. As a result, this paper captures the trade-off between the volume of data transmitted and the resulting increase in transmission delay while optimizing the task performance at the destination.
Furthermore, [1] established the optimal threshold-based waiting-time rule for the time-invariant packet length $l=1$ and showed that the freshness of the data packets to be transmitted is a function only of the network delay state in the previous epoch. The present paper extends these results by generalizing the waiting-time rule to arbitrary time-invariant packet lengths $l$ (not limited to $l=1$) and by deriving an exact closed-form expression for the freshness of the data packets to be transmitted. Finally, this paper also investigates scenarios where the packet length can be adjusted over time based on the state of the scheduler in Section IV.
@@ II System Model
We consider a time-slotted remote inference system, which comprises a data source, a transmitter, and a receiver, as illustrated in Fig. 1. On the transmitter side, the source signal $V_{t}$ is regularly sampled in each time slot $t$. The samples are stored in a buffer containing the most recent $B$ samples $\big(V_{t},V_{t-1},\ldots,V_{t-B+1}\big)$, waiting to be sent to the receiver. On the receiver side, a predictor, e.g., a pre-trained neural network, is used to infer the real-time value of the target signal $Y_{t}$ of interest. In practical time-series forecasting algorithms, the predictor needs to use multiple consecutive samples, e.g., $X_{t-b}^{l}=(V_{t-b},V_{t-b-1},\ldots,V_{t-b-l+1})$, to infer $Y_{t}$. We will design a transmission scheduler that determines (i) when to send a group of consecutive samples over the communication network to the receiver and (ii) which group of consecutive samples in the buffer to send. This medium access model is called “selection-from-buffer,” which was proposed recently in [6, 7]. Such remote inference systems are crucial for numerous real-time applications, including sensor networks, airplane/vehicular control, robotics networks, and digital twins, where timely and accurate remote monitoring and control are essential for maintaining system stability and performance.
@@
We assume that the system starts operation at time slot $t=0$. In the selection-from-buffer model, a new sample of the source signal $V_{t}\in\mathcal{V}$ is added to the buffer at the beginning of each time slot $t$, and the oldest sample is discarded. The buffer, with size $B$, is assumed to be initially full, storing the samples $\big(V_{0},V_{-1},\ldots,V_{-B+1}\big)$. Thus, at any time slot $t$, the buffer contains the most recent $B$ samples $\big(V_{t},V_{t-1},\ldots,V_{t-B+1}\big)$. The transmitter sends status update packets to the receiver one by one. Let $S_{i}$ and $D_{i}$ represent the transmission start time and delivery time of the $i$-th transmitted packet, respectively, such that $S_{i}<D_{i}$. At time slot $t=S_{i}$, the scheduler selects a group of consecutive samples from the buffer and forms the packet $X_{S_{i}-b_{i}}^{l_{i}}=\big(V_{S_{i}-b_{i}},V_{S_{i}-b_{i}-1},\ldots,V_{S_{i}-b_{i}-l_{i}+1}\big)\in\mathcal{V}^{l_{i}}$. Here, $l_{i}\in\{1,\ldots,B\}$ specifies the sequence length of the $i$-th packet $X_{S_{i}-b_{i}}^{l_{i}}$, while $b_{i}$ denotes the relative position of the sample $V_{S_{i}-b_{i}}$ in the buffer. The latter parameter $b_{i}$ controls the freshness of the samples included in the packet. When $B=1$, the selection-from-buffer model reduces to the generative-at-will model which has been studied extensively in previous works [37, 18, 38, 39, 19, 27, 28, 29]. Hence, the packet $X_{S_{i}-b_{i}}^{l_{i}}$ is submitted to the communication network at time slot $S_{i}$ and delivered to the receiver at time slot $D_{i}$. Upon delivery, the receiver sends an acknowledgment (ACK) back to the transmitter, which is received at time slot $A_{i}$.
We assume that the scheduler always waits for feedback before transmitting a new data packet. In other words, the scheduler remains silent during the time slots between $D_{i}$ and $A_{i}$ such that $S_{i}<D_{i}<A_{i}\leq S_{i+1}$ for all $i$. This assumption clearly restricts the throughput, the maximization of which has been the main objective in conventional user-centric, pipelined communication systems. However, under the concept of goal-oriented communication, which has been introduced to address the scalability problem in massive machine-type networks, the main objective is not throughput maximization but ensuring the effective execution of machine tasks at the destination while operating under limited communication resources. From this perspective, allowing multiple packets to be simultaneously in flight may increase network load without yielding proportional improvements in task performance, particularly in dense environments where congestion and contention are the primary concerns. Consequently, the one-packet-in-flight (stop-and-wait) assumption serves as a scenario-driven modeling choice that captures conservative communication behavior suitable for the class of systems considered in this work.